Project

General

Profile

NDOS CommissionMtng 20121008
10/08/2012
Attending: Mathew Muether, Gavin Davies, Zukai Wang, Jan Zernstein, Xuebing Bu, Peter Shanahan, Venktesh Singh

Weekly Topics:
Update on FarDet snout test stand - Junting will be coming to Fermilab for 1 month to work on this following the collaboration meeting

-Zukai NDOS noise study -
1 channel on each APD
3 coating type
parylene 3-2 (14,5) (X)
2-2 (22, 23) (X2)
3-3 (42,45) (X3)
Andrew is back tomorrow. Zukai prefers to wait until then.

-DCM, FEB firmware upgrades and new IOC -DCM Firmware rollout. This wil be done this

-DCS errors on the Labview environmental monitor - Athans and I will investigate cabling now Wednesday.

-Slow DB - Experts have been notified

-RC Server death - DAQ application manager could send alarms and shutdown datalogger. Experts have been notified

UPDATES:
-Detector Update - See above
-DataCheck Update - No update (DB migration holdup continues)
Jan - DCSReport table is empty on the if machine.

-Installation Plans for the week - None
-DAQ Update and Plans for the week - See Above
-DCS Update and Plans for the week - None